OpenStudy (anonymous):

steps to find prime factorization of 1260

OpenStudy (anonymous):

smallest prime is 2 so start by dividing the number by 2

OpenStudy (anonymous):

1260=2*630 630 is divisible by 2 so continue dividing by 2 1260=2*630=2*2*315 315 is not divisible by 2 so go for dividing by 3 which is a prime number greater than 2 and ...

OpenStudy (mayankdevnani):

 \[2\times2\times3\times3\times7\times5=1260\]